# Condensate Size Control by Charge Asymmetry

### v1

@@ B.2 Numerical minimization method
We designed an iterative scheme based on the self-consistent equations above, where we update all values of fields and variables based on their current approximated values, so the iteration converges to the free energy minimum. In this scheme, we first calculate $Q_{i}$ using Eq. (S13), and $\phi_{i}(\beta,\mathrm{r})$ using Eq. (S15). We then calculate $\psi(\beta,\mathrm{r})$ via
@@
where $k$ is the wavenumber of the Fourier transform $\mathcal{F}$ and its inverse $\mathcal{F}^{-1}$. Next, we calculate $\xi(\beta,\mathrm{r})$ using Eq. (B.1) and the incompressibility condition given by Eq. (S21),
@@
We now use Eq. (B.1) to update the ideal part of the new field, $\omega_{i}^{\mathrm{ideal}}(\beta,\mathrm{r})$, via
@@
where $A_{\omega}$ is an empirical acceptance rate, typically around $0.001$. We next use Eq. (B.1) to update ${J_{\beta}}$. More specifically, using
@@
Here, $A_{J}$ is another empirical acceptance rate, which is typically set to 0.001. Afterwards, we shift $J^{\mathrm{new}}_{\beta}$ to satisfy Eq. (S22) which reads $\sum_{\beta}J^{\mathrm{new}}_{\beta}=1$. Next, we update $\zeta(\beta)$ using Eq. (S20),
@@
where $A_{\zeta}$ is the third acceptance rate, which is set to $10$. Finally, in 1D, we update $L_{\beta}$ using Eq. (S27), via
$L^{\mathrm{new}}_{\beta}$ $=$ $L_{\beta}-A_{L}\times\left(-{J_{\beta}}\frac{2}{{L_{\beta}}}\frac{1}{{L_{\beta}}}\int_{0}^{L_{\beta}}\Bigg{[}-\frac{1}{2}\sum_{i}{\kappa}_{i}\phi_{i}(\beta,x)\partial_{x}^{2}\phi_{i}(\beta,x)+\frac{{v}}{8\pi}\psi(\beta,x)\partial_{x}^{2}\psi(\beta,x)\bigg{]}\mathrm{d}x\right)\;,$ (S36)
where $A_{L}$ is an empirical acceptance rate around $10$. We use this iterative scheme to find the equilibrium states, where each phase is described by fields $\phi_{i}(\mathrm{r})$ and $\psi(\mathrm{r})$ together with optimized sizes $L_{\beta}$. To produce the results of the main text, we use ${N_{\mathrm{c}}}=5$, assume $M=2$, and discretize fields at $128$ points along the one dimension. We have checked the correctness at selected parameter values by using $M=8$; All convergent results are consistent with those using $M=2$.
@@ B.3 Linear stability analysis
We here present the linear stability analysis for the uniform state $\phi_{i}(x)=\bar{\phi}_{i}$. We perturb $\bar{\phi}_{i}$ by a small value, setting $\phi_{i}(x)=\bar{\phi}+\epsilon_{i}\cos(qx)$, where $\epsilon_{i}$ is the perturbation amplitude and $q$ is the associated wave number. By evaluating Eq. (S7), taking the second-order derivative of $\hat{f}$ with respect to $\epsilon_{i}$, and then taking the limit $\epsilon_{i}\rightarrow 0$, we obtain the $4\times 4$ Hessian matrix
@@
For the homogeneous state to be stable, all eigenvalues $\lambda_{i}$ of the Hessian matrix must be positive for all $q$. If the homogeneous state is unstable, we numerically calculate the eigenvalues and identify the lowest $\lambda_{i}$ as a function of $q$. The length scale of this most unstable mode is given by $2\pi/q$ and shown in Fig. 3d of the main text. We also show the most unstable length scale for small charge asymmetry in Fig. S5, indicating that this length strongly depends on $\chi$, while the charge asymmetry has a weak effect, in contrast to the results for the equilibrium length scale.
@@
Figure S5: The most unstable length scale predicted from linear stability analysis. Parameters are the same as in Fig. 3(d) of the main text.
B.4 Additional result figures
@@
Figure S6 shows phase diagram of states, pattern periods, amplitudes, and electrostatic energy for $\chi=-4$ and $\chi=-7$, analogous to Fig. 3b in the main text. Figures S7 and S8 provide additional details for states depicted in Fig. 2 of the main text. Figure S9 shows a typical profile near the continuous transition. This transition is demonstrated by the amplitude of the polycation volume fraction, as shown in Fig. S10. The different contributions to the free energy are shown in Fig. S11. The last two columns, in particular, demonstrate that the interfacial energy $F_{\mathrm{int}}$ is equal to the electrostatic energy $F_{\mathrm{el}}$.
@@
We also obtain the patterned phase by solving the two-dimensional mean-field theory equations. We show the profiles of volume fractions in Fig. S12 and charge densities in Fig. S13, corresponding to the same states as in Fig. 2 of the main text. The 2D disk patterns qualitatively agree with the 1D results; notably, we also observe a dip inside the droplet at $z_{\mathrm{C}}=0.4$.
Figure S6: Results analogous to Fig. 3 in the main text for $\chi=-4$ (upper row) and $\chi=-7$ (lower row).
@@
Figure S7: Additional details for the state for weak charge asymmetry shown in Fig. 2a in the main text. Parameters are $z_{\mathrm{C}}=0.4$, $z_{\mathrm{A}}=-0.2$, $\chi=-5$, and $\kappa=10$. (a) Volume fractions $\phi_{i}$. (b) Electrostatic potential $\psi$. (c) Local electrostatic energy density $\rho\psi$. (d) Charge density $\rho_{i}=z_{i}\phi_{i}$ for each species. (e) Total charge density $\rho=\sum_{i}\rho_{i}$. (f) Absolute value of total charge density.
@@
Figure S8: Additional details for the state for weak charge asymmetry shown in Fig. 2b in the main text. Parameters are $z_{\mathrm{C}}=0.8$, $z_{\mathrm{A}}=-0.2$, $\chi=-5$, and $\kappa=10$. (a) Volume fractions $\phi_{i}$. (b) Electrostatic potential $\psi$. (c) Local electrostatic energy density $\rho\psi$. (d) Charge density $\rho_{i}=z_{i}\phi_{i}$ for each species. (e) Total charge density $\rho=\sum_{i}\rho_{i}$. (f) Absolute value of total charge density.
@@
Figure S9: Typical profile close to continuous transition. Parameters are $z_{\mathrm{C}}=0.8$, $z_{\mathrm{A}}=-0.2$, $\chi=-4.5$ (the critical point is at $\chi^{*}\approx-4.46$), and $\kappa=10$. (a) Volume fractions $\phi_{i}$. (b) Electrostatic potential $\psi$. (c) Local electrostatic energy density $\rho\psi$. (d) Charge density $\rho_{i}=z_{i}\phi_{i}$ for each species. (e) Total charge density $\rho=\sum_{i}\rho_{i}$. (f) Absolute value of total charge density.
@@
Figure S10: Amplitude decreases as $\chi$ increases. (a) The amplitude of polycation profile $\phi_{C}$ decreases as $\chi$ increases. The circle indicates the critical point at $\chi=\chi^{*}$ where $\chi^{*}\approx-4.46$ for $z_{\mathrm{C}}=0.8$ and $\chi^{*}\approx-4.12$ for $z_{\mathrm{C}}=0.6$. (b) Scaling of the amplitude as a function of the relative interaction strength $\chi^{*}-\chi$. The exponent $\frac{1}{2}$ is consistent with the critical exponent in mean-field theory, indicating that the transition is continuous. Parameters are $z_{\mathrm{A}}=-0.2$ and $\kappa=10$. Remaining parameters are the same as in Fig. 2 in the main text.
@@
Figure S11: Different contributions to the free energy. Total energy density $F$, and its three contributions $F_{\mathrm{FH}}$, $F_{\mathrm{int}}$, and $F_{\mathrm{el}}$, normalized to the size $L$ as a function of the charges $z_{C}$ and $z_{A}$ of the polyions for various interaction strengths $\chi$. To highlight the details $F$ and $F_{\mathrm{FH}}$ have been shifted by the free energy $F(\phi_{i}(x)=\bar{\phi}_{i}$) of the respective homogeneous states. Consistent with Eq. (S26), $F_{\mathrm{int}}$ and $F_{\mathrm{el}}$ are always identical.
@@
Figure S12: Two-dimensional profiles of the volume fractions $\phi_{i}(x,y)$ predicted by the field theory. The parameters are the same as in Fig. 2 of the main text.
@@
Figure S13: Two-dimensional profiles of the charge densities, $\rho_{i}(x,y)=z_{i}\phi_{i}(x,y)$, predicted by the field theory. The last column shows the total charge density, $\sum_{i}\rho_{i}(x,y)$.

### v2

@@ B.2 Numerical minimization method
We designed an iterative scheme based on the self-consistent equations above, where we update all values of fields and variables based on their current approximated values, so the iteration converges to the free energy minimum. To improve convergence, we extensively use a simple mixing formula to improve numerical stability, i.e., for a generic variable $X$, we determine its value $X^{\text{new}}$ at the next iteration step as $X^{\mathrm{new}}=X^{\text{old}}+A_{X}(X^{\text{ideal}}-X^{\text{old}})$, where $X^{\mathrm{old}}$ is the current value and $X^{\mathrm{ideal}}$ is the value suggested by the iteration scheme. Here, $A_{X}$ is an empirically determined mixing rate.
In our numerical scheme, we first calculate $Q_{i}$ using Eq. (S13), and $\phi_{i}(\beta,\mathrm{r})$ using Eq. (S15). We then calculate $\psi(\beta,\mathrm{r})$ via
@@
where $k$ is the wavenumber of the Fourier transform $\mathcal{F}$ and its inverse $\mathcal{F}^{-1}$. Next, we calculate $\xi(\beta,\mathrm{r})$ using Eq. (S17) and the incompressibility condition given by Eq. (S21),
@@
We now use Eq. (S17) to update the ideal part of the new field, $\omega_{i}^{\mathrm{ideal}}(\beta,\mathrm{r})$, via
@@
where $A_{\omega}$ is an empirical mixing rate, typically around $0.001$. In fact, Eq. (S31) is a simple mixing strategy together with a low pass filter, where the iteration is faster for low frequencies while it is slower slower for high frequencies to make the self-consistent iteration scheme more stable [53, 54]. This approach empirically improved the convergence. We next use Eq. (B.1) to update ${J_{\beta}}$. More specifically, using
@@
Here, $A_{J}$ is another empirical mixing rate, which is typically set to 0.001. Afterwards, we shift $J^{\mathrm{new}}_{\beta}$ to satisfy Eq. (S22) which reads $\sum_{\beta}J^{\mathrm{new}}_{\beta}=1$. Next, we update $\zeta(\beta)$ using Eq. (S20),
@@
where $A_{\zeta}$ is the third mixing rate, which is set to $10$. Finally, in 1D, we update $L_{\beta}$ using Eq. (S27), via
$L^{\mathrm{new}}_{\beta}$ $=$ $L_{\beta}+A_{L}{J_{\beta}}\frac{2}{{L_{\beta}}}\frac{1}{{L_{\beta}}}\int_{0}^{L_{\beta}}\bigg{[}-\frac{1}{2}\sum_{i}{\kappa}_{i}\phi_{i}(\beta,x)\partial_{x}^{2}\phi_{i}(\beta,x)+\frac{{v}}{8\pi}\psi(\beta,x)\partial_{x}^{2}\psi(\beta,x)\bigg{]}\mathrm{d}x\;,$ (S36)
where $A_{L}$ is an empirical mixing rate around $10$. We use our iterative scheme to find equilibrium states where each phase is described by fields $\phi_{i}(\mathrm{r})$ and $\psi(\mathrm{r})$ together with optimized sizes $L_{\beta}$. To produce the results of the main text, we use ${N_{\mathrm{c}}}=5$, assume $M=2$, and discretize fields at $128$ points along the single dimension. We have checked the correctness at selected parameter values using $M=8$ phases; All convergent results are consistent with those using $M=2$.
@@ B.3 Linear stability analysis
We here present the linear stability analysis of the field theory around the uniform state $\phi_{i}(x)=\bar{\phi}_{i}$. We perturb $\bar{\phi}_{i}$ by a small value, setting $\phi_{i}(x)=\bar{\phi}+\epsilon_{i}\cos(qx)$, where $\epsilon_{i}$ is the perturbation amplitude and $q$ is the associated wave number. By evaluating Eq. (S7), taking the second-order derivative of $\hat{f}$ with respect to $\epsilon_{i}$, and then taking the limit $\epsilon_{i}\rightarrow 0$, we obtain the $4\times 4$ Hessian matrix
@@
For the homogeneous state to be stable, all eigenvalues $\lambda_{i}$ of the Hessian matrix must be positive for all $q$, which we check numerically. If the homogeneous state is unstable, we numerically calculate the eigenvalues as a function of $q$ and determine the minimal value at $q=q_{\mathrm{min}}$. The length scale associated with this most unstable mode is $2\pi/q_{\mathrm{min}}$ and shown in Fig. 4 of the main text. It provides an estimate for the length scale originating when an unstable homogeneous system is prepared, and it sometimes provides a reasonable estimate for the final pattern length scale. To test this, we show the most unstable length scale for small charge asymmetry in Fig. S18, indicating that this length strongly depends on $\chi$, while the charge asymmetry has a weak effect, in contrast to the results for the equilibrium length scale presented in the main text.
@@
Figure S18: The most unstable length scale $L$ predicted from linear stability analysis as a function of the interaction parameter $\chi$. Parameters are the same as in Fig. 4 of the main text.
B.4 Quantification of ion concentration
One of the main messages of the main text is that ions are unevenly distributed in the patterned phase, with a higher concentration of ions outside the droplet compared to the interior. We here quantify this effect by calculating the average fraction of ions inside droplets and outside droplets separately. Fig. S19 shows that ions are enriched outside droplets (right column) while they are depleted inside droplets (left columns) in the patterned phase. This disparity arises because the short-range attraction between polymers effectively expels ions from the droplet. The strength of this expulsion increases with stronger polymer attractions, (larger $-\chi$) as illustrated in Fig. S20. However, the expulsion weakens as charge asymmetry increases; see Fig. S19.
Figure S19: The relative mean volume fraction of ions, $\phi_{\mathrm{ion}}/\bar{\phi}_{\mathrm{ion}}$, for (a) inside and (b) outside the droplet as a function of the charge numbers of the polymers. Here, $\phi_{\mathrm{ion}}$ is the mean volume fraction of all small ions ($e^{+}$ and $e^{-}$) inside or outside the droplet, and $\bar{\phi}_{\mathrm{ion}}=\bar{\phi}_{e^{+}}+\bar{\phi}_{e^{-}}$. The inside of the droplet is defined as the region where $\phi_{\mathrm{S}}<\frac{1}{2}(\phi_{\mathrm{S}}^{\mathrm{min}}+\phi_{\mathrm{S}}^{\mathrm{max}})$, with $\phi_{\mathrm{S}}^{\mathrm{min}}$ and $\phi_{\mathrm{S}}^{\mathrm{max}}$ representing the minimum and maximum values of the solvent volume fraction, respectively. Conversely, regions where $\phi_{\mathrm{S}}\geq\frac{1}{2}(\phi_{\mathrm{S}}^{\mathrm{min}}+\phi_{\mathrm{S}}^{\mathrm{max}})$ are considered outside the droplet.
Figure S20: The relative mean volume fraction of ions, $\phi_{\mathrm{ion}}/\bar{\phi}_{\mathrm{ion}}$, as a function of the interaction strength $\chi$ for various charge numbers of the polymers. A stronger short-range attraction (more negative $\chi$) results in a greater expulsion of ions from the droplet.
B.5 Additional result figures
@@
Figure S21 shows phase diagram of states, pattern periods, amplitudes, and electrostatic energy for $\chi=-4$ and $\chi=-7$, analogous to Fig. 3(b) in the main text. Figures S22 and S23 provide additional details for states depicted in Fig. 2 of the main text. Figure S24 shows a typical profile near the continuous transition. The nature of this transition can be investigate from the dependence of the amplitude of the volume fraction of the negatively-charged polymer $P_{-}$ as a function of $\chi$; see Fig. S25. The different contributions to the free energy are shown in Fig. S26. The last two columns, in particular, demonstrate that the interfacial energy $F_{\mathrm{int}}$ is equal to the electrostatic energy $F_{\mathrm{el}}$.
@@
To check whether dimensionality matters, we also obtain the patterned phase by solving the two-dimensional mean-field theory equations. We show the profiles of volume fractions in Fig. S27 and charge densities in Fig. S28, corresponding to the same states as in Fig. 2 of the main text. The 2D disk patterns qualitatively agree with the 1D results; Notably, we also observe a dip inside the droplet at $z_{P^{+}}=0.4$. Given the low total volume fractions of polymers, $\phi_{P^{+}}+\phi_{P^{-}}=0.2$, we expect spheres, rather than lamellar or cylindrical structures, will form in 3D within our parameter regime.
Figure S21: Results analogous to Fig. 3 in the main text for $\chi=-4$ (upper row) and $\chi=-7$ (lower row). (a) Phase diagram as a function of the charge numbers $z_{P^{+}}$ and $z_{P^{-}}$ of the polymers revealing parameter regions with the coexistence of two homogeneous phases (H+H), patterned phases (P), a single homogeneous phase (H). (b) Period $L$ in the patterned phase corresponding to panel a. (c) Amplitude $A_{P^{+}}=\max(\phi_{P^{+}})-\min(\phi_{P^{+}})$ corresponding to panel a. (d) Electrostatic energy $F_{\mathrm{el}}$ corresponding to panel a.
@@
Figure S22: Additional details for the state for weak charge asymmetry shown in Fig. 2a in the main text. Parameters are $z_{P^{+}}=0.4$, $z_{P^{-}}=-0.2$, $\chi=-5$, and $\kappa=10$. (a) Volume fractions $\phi_{i}$ as a function of position $x$. (b) Electrostatic potential $\psi(x)$. (c) Local electrostatic energy density $\rho\psi(x)$. (d) Charge density $\rho_{i}(x)=z_{i}\phi_{i}(x)$ for each species. (e) Total charge density $\rho(x)=\sum_{i}\rho_{i}(x)$. (f) Absolute value of total charge density $|\rho(x)|$.
@@
Figure S23: Additional details for the state for weak charge asymmetry shown in Fig. 2b in the main text. Parameters are $z_{P^{+}}=0.8$, $z_{P^{-}}=-0.2$, $\chi=-5$, and $\kappa=10$. (a) Volume fractions $\phi_{i}$ as a function of position $x$. (b) Electrostatic potential $\psi(x)$. (c) Local electrostatic energy density $\rho\psi(x)$. (d) Charge density $\rho_{i}(x)=z_{i}\phi_{i}(x)$ for each species. (e) Total charge density $\rho(x)=\sum_{i}\rho_{i}(x)$. (f) Absolute value of total charge density $|\rho(x)|$.
@@
Figure S24: Typical profile close to continuous transition. Parameters are $z_{P^{+}}=0.8$, $z_{P^{-}}=-0.2$, $\chi=-4.5$ (the critical point is at $\chi^{*}\approx-4.46$), and $\kappa=10$. (a) Volume fractions $\phi_{i}$ as a function of position $x$. (b) Electrostatic potential $\psi(x)$. (c) Local electrostatic energy density $\rho\psi(x)$. (d) Charge density $\rho_{i}(x)=z_{i}\phi_{i}(x)$ for each species. (e) Total charge density $\rho(x)=\sum_{i}\rho_{i}(x)$. (f) Absolute value of total charge density $|\rho(x)|$.
@@
Figure S25: Amplitude decreases as $\chi$ increases. (a) The amplitude of fraction $\phi_{P^{+}}$ of the positively-charged polymer decreases as the interaction parameter $\chi$ increases. The circle indicates the critical point at $\chi=\chi^{*}$ where $\chi^{*}\approx-4.46$ for $z_{P^{+}}=0.8$ and $\chi^{*}\approx-4.12$ for $z_{P^{+}}=0.6$. (b) Rescaled data of panel a, now shown as a function of the relative interaction strength $\chi^{*}-\chi$. The exponent $\frac{1}{2}$ is consistent with the critical exponent in mean-field theory, indicating that the transition is continuous. Parameters are $z_{P^{-}}=-0.2$ and $\kappa=10$. Remaining parameters are the same as in Fig. 2 in the main text.
@@
Figure S26: Different contributions to the free energy. Total energy density $F$, and its three contributions $F_{\mathrm{FH}}$, $F_{\mathrm{int}}$, and $F_{\mathrm{el}}$, normalized to the size $L$ as a function of the charge numbers $z_{P^{+}}$ and $z_{P^{-}}$ of the polymers for various interaction strengths $\chi$ (across rows). To highlight the details, $F$ and $F_{\mathrm{FH}}$ have been shifted by the free energy $F(\phi_{i}(x)=\bar{\phi}_{i}$) of the respective homogeneous states. Consistent with Eq. (S26), $F_{\mathrm{int}}$ and $F_{\mathrm{el}}$ are always identical.
@@
Figure S27: Two-dimensional profiles the volume fractions $\phi_{i}(x,y)$ of all species (across columns) predicted by the field theory for two different charge numbers $z_{P^{+}}$ of the positively-charged polymers (across rows). The remaining parameters are the same as in Fig. 2 of the main text.
@@
Figure S28: Two-dimensional profiles of charge densities, $\rho_{i}(x,y)=z_{i}\phi_{i}(x,y)$, of charged species (across columns) predicted by the field theory for two charge numbers $z_{P^{+}}$ of the positively-charged polymers (across rows). Last column shows the total charge density $\sum_{i}\rho_{i}(x,y)$.